A Bibliography of Publications of Radha Jagadeesan

Radha Jagadeesan

Dept. of Mathematical and Computer Sciences

Loyola University—Lake Shore Campus

6525 N. Sheridan Road

Chicago, IL 60626 USA

Tel: +1 773 508 3526 FAX: +1 773 508 3514

E-mail: radha@cs.luc.edu (Internet)

13 October 2017 Version 1.10

Abstract

This bibliography records publications of Radha Jagadeesan.

Title word cross-reference

+ [CJJ^{+}98c]. L [Jag89, Jag90]. λ [JJPR06, JJPR08]. RBAC [JJPR06].

-domains [Jag89, Jag90]. -RBAC [JJPR08].

17th [Pat90].

4th [AN95].

5th [Mai90].

7th [MS96].

’95 [AN95]. ’96 [MS96].

Abstract [JP92, AJ92a, AMJ94, JPP91, SJ05]. Abstraction [GHJ01, AMJ94, AJM00, GJ02]. Abstraction-based [GHJ01]. access [JJPR06, JJPR08]. Albuquerque [ACM92]. Algebraic [AN95, CJJ^{+}98a, CJJ^{+}98c]. AMAST [AN95]. Analysis [JJ95]. Annual [ACM92]. API [CJJ^{+}98a, CJJ^{+}98c]. Approximate [GJP06]. Approximating [DJGP00, DGJP03]. April [Mai90]. aspects [JJR06]. August [MS96]. Automata [GHJ97, Pat90, GJS96a]. Automatic [GJ02].

based [GHJ01, JJPR06, JJPR08]. bisimulation [DGJP02, DGJP10].

Calculus [JP90, JJ95]. Canada [AN95]. case [CJJ^{+}98b]. Causality [JJ95]. cc [GJS96a]. change [GJS98]. checking [GHJ01, GJ02]. clustered [SJ05]. colloquium [Pat90]. complete [DGJP02, DGJP10]. completeness [AJ92a, AJ94a, AJ95]. Computer [AJL07]. Computing [GJS98]. CONCUR [MS96]. Concurrency [JJ95, MS96]. Concurrent [SJG94a, SJG94b, SJ05, GJS96b, GJS96c, GJS97, GJS02, JNS05, SJG96]. Conference [ACM92, MS96, AN95, AJL07, Mai90]. Constraint [GJSB95, SJG94a, SJG94b, GJS96b, GJS96c, GJS97, GJS02, SJG96]. continuous [GJS98]. control [JJPR06, JJPR08].

Dataflow [JJ95]. Default [SJG96]. Design [CJJ^{+}98a]. Domain [JP90, HJS04]. Domain-theoretic [JP90]. domains [Jag89, Jag90].

England [Pat90]. equation [HJS04]. Events [CJJ^{+}98a, CJJ^{+}98c]. extended [AJ92a, AMJ94, SJ05].

first [JPP91]. Foundations [AJ92b, AJ94b, SJG94a, Mai90]. Full [AMJ94, AJM00, AJ92a, AJ94a, AJ95]. fully [JPP91]. fully-abstract [JPP91]. Functional [JP92, JPP91].

game [AJ03, AJ05]. Games [AJ92a, AJ94a, AJ95]. generalized [GJ02]. generic [AJ03, AJ05]. Geometry [AJ92b, AJ94b].

Higher [JP90, JP92]. Higher-Order [JP92, JP90]. Hybrid [GJSB95, GJS96a].

Implementation [CJJ^{+}98a]. Interaction [AJ92b, AJ94b]. International [MS96, AN95, Mai90, Pat90]. interpretation [JNS05]. intuitionistic [JNS05]. issue [AJL07]. Italy [MS96].

January [ACM92]. Java [CJJ^{+}98b]. July [AN95, Pat90]. labeled [DGJP99, DJGP00]. labelled [DGJP03, DGJP04]. Language [JP92, JPP91]. Languages [ACM92, GJSB95, SJG94b, Pat90]. linear [AJ92a, AJ94a, AJ95]. Logic [AJL07, AJ92a, AJ94a, AJ95, JPP91, JP92, JNS05]. Lossless [Jag89, Jag90]. Louisiana [Mai90].

March [Mai90]. Markov [DGJP99, DJGP00, DGJP03, DGJP04]. Mathematical [Mai90]. methodology [AN95]. Metrics [DGJP99, DGJP04]. Mexico [ACM92]. modal [GHJ01]. Model [JP90, GHJ01, GJ02]. Models [GJS96b]. Montreal [AN95]. MR2375571 [JJPR08]. multiplicative [AJ92a, AJ94a, AJ95].

Nineteenth [ACM92].

Objects [CJJ^{+}98b]. Order [JP92, JP90, JPP91]. Orleans [Mai90].

papers [ACM92, AJL07]. parametric [JJR06]. partial [HJS04]. PCF [AMJ94, AJM00]. pCTL [DGJP10]. PCTL* [DGJP02]. Pi [JJ95]. Pi-Calculus [JJ95]. Pisa [MS96]. polymorphism [AJ03, AJ05, JJR06]. Powerdomains [Jag89, Jag90]. Preface [AJL07]. presented [ACM92]. Principles [ACM92]. Probabilistic [GJS97, GJP06]. proceedings [AN95, Mai90, MS96, Pat90]. Process [CJJ^{+}98a, JP90, CJJ^{+}98c]. Process-algebraic [CJJ^{+}98a, CJJ^{+}98c]. Processes [CJJ^{+}98b, DJGP00, DGJP03, DGJP04, GJP06]. program [GJS96a]. Programming [ACM92, GJSB95, SJG94a, SJG94b, GJS96b, GJS96c, GJS97, GJS02, JJPR06, JJPR08, Mai90, Pat90, SJG96, SJ05].

RBAC [JJPR08]. real [GJP06]. real-time [GJP06]. reasoning [GJP06]. record [ACM92]. refinement [HJS04]. reprint [JJPR08]. Robust [GHJ97]. role [JJPR06, JJPR08]. role-based [JJPR06, JJPR08].

Science [AJL07]. selected [AJL07]. Semantics [JP92, AJ03, AJ05, CJJ^{+}98c, JPP91, Mai90]. SIGACT [ACM92]. SIGPLAN [ACM92]. SIGPLAN-SIGACT [ACM92]. software [AN95]. sound [DGJP02, DGJP10]. Special [AJL07]. study [CJJ^{+}98b]. Symposium [ACM92]. systems [DGJP99, GHJ01, HJS04, JNS05].

technology [AN95]. telecommunication [CJJ^{+}98b]. Testing [JNS05]. theoretic [JP90]. theory [MS96]. Threads [CJJ^{+}98a, CJJ^{+}98c]. time [GJP06]. Timed [GHJ97, SJG94a, SJG94b, SJG96]. transition [GHJ01]. Triveni [CJJ^{+}98a, CJJ^{+}98b, CJJ^{+}98c]. True [JJ95]. Truly [GJS96c, GJS02]. Tulane [Mai90]. Typed [JJR06].

University [Mai90, Pat90]. USA [Mai90]. using [GHJ01, GJ02].

variables [JPP91, JP92]. verification [GJS96a].

Warwick  [Pat90].  Weak

[DGJP02, DGJP10].

References

ACM:1992:CRN

[ACM92] ACM, editor. Conference record of the Nineteenth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Albuquerque, New Mexico, January 19–22, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. ACM order number 54990.

Abramsky:1992:GFC

[AJ92a] Samson Abramsky and Radha Jagadeesan. Games and full completeness for multiplicative linear logic (extended abstract). In Foundations of software technology and theoretical computer science (New Delhi, 1992), volume 652 of Lecture Notes in Comput. Sci., pages 291– 301. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992.

Abramsky:1992:NFG

[AJ92b] Samson Abramsky and Radha Jagadeesan. New foundations for the geometry of interaction. In Proceedings 7th Annual IEEE Symp. on Logic in Computer Science, LICS’92, Santa Cruz, CA, USA, 22–25 June 1992, pages 211–222. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992.

Abramsky:1994:GFC

[AJ94a] Samson Abramsky and Radha Jagadeesan. Games and full completeness for multiplicative linear logic. Journal of Symbolic Logic, 59(2): 543–574, 1994. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

Abramsky:1994:NFG

[AJ94b] Samson Abramsky and Radha Jagadeesan. New foundations for the geometry of interaction. Information and Computation, 111(1):53– 119, May 15, 1994. CODEN IN-FCEC. ISSN 0890-5401 (print), 1090-2651 (electronic). 1992 IEEE Symposium on Logic in Computer Science.

Abramsky:1995:GFC

[AJ95] S. Abramsky and R. Jagadeesan.

Games and full completeness for multiplicative linear logic. Journal of Symbolic Logic, 1995. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

Abramsky:2003:GSG

[AJ03] Samson Abramsky and Radha Jagadeesan. A game semantics for generic polymorphism. In Foundations of software science and computation structures, volume 2620 of Lecture Notes in Comput. Sci., pages 1–22. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Abramsky:2005:GSG

[AJ05] Samson Abramsky and Radha Jagadeesan. A game semantics for generic polymorphism. Ann. Pure Appl. Logic, 133(1-3):3–37, 2005. CODEN APALD7. ISSN 0168-0072.

Alur:2007:PSI

[AJL07] Rajeev Alur, Radha Jagadeesan, and Leonid Libkin. Preface [Special issue: selected papers of the conference “Logic in Computer Science 2006”]. Log. Methods Comput. Sci., ??(??):[front matter], 2007. ISSN 1860-5974. Special issue: Conference “Logic in Computer Science 2006”, held in Seattle, WA, August 12–15, 2006.

Abramsky:2000:FAP

[AJM00] Samson Abramsky, Radha Jagadeesan, and Pasquale Malacaria. Full abstraction for PCF. Information and Computation, 163(2):409– 470, 2000. CODEN INFCEC. ISSN 0890-5401 (print), 1090-2651 (electronic).

Abramsky:1994:FAP

[AMJ94] Samson Abramsky, Pasquale Malacaria, and Radha Jagadeesan. Full abstraction for PCF (extended abstract). Lecture Notes in Computer Science, 789:1–15, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagar:1995:AMS

[AN95] Vangalur S. Alagar and M. Nivat, editors. Algebraic methodology and software technology: 4th international conference, AMAST ’95, Montreal, Canada, July 3–7, 1995: proceedings, volume 936 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60043-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 1995.

Colby:1998:DIT

[CJJ^{+}98a] C. Colby, L. J. Jagadeesan, R. Jagadeesan, K. Läufer, and C. Puchol. Design and implementation of Triveni: A process-algebraic API for threads + events. In Proceedings of the 1998 IEEE International Conference on Computer Languages. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. To appear.

Colby:1998:OPT

[CJJ^{+}98b] C. Colby, L. J. Jagadeesan, R. Jagadeesan, K. Läufer, and C. Puchol. Objects and processes in Triveni: A telecommunication case study in Java. In Proceedings of the 1998 Usenix Conference on Object Oriented Technologies and Systems, page ?? USENIX, Berkeley, CA, USA, 1998. To appear.

Colby:1998:STP

[CJJ^{+}98c] Christopher Colby, Lalita Jategaonkar, Radha Jagadeesan, Konstantin Läufer, and Carlos Puchol. The semantics of Triveni: a processalgebraic API for threads + events. In US–Brazil Joint Workshop on the Formal Foundations of Software Systems (Rio de Janeiro, 1997/ New Orleans, LA, 1997), volume 14 of Electron. Notes Theor. Comput. Sci., pages 1–27. Elsevier, Amsterdam, The Netherlands, 1998.

Desharnais:1999:MLM

[DGJP99] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Metrics for labeled Markov systems. In CONCUR’99: concurrency theory (Eindhoven), volume 1664 of Lecture Notes in Comput. Sci., pages 258–273. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999.

Desharnais:2002:WBS

[DGJP02] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Weak bisimulation is sound and complete for PCTL*. In CONCUR 2002—concurrency theory, volume 2421 of Lecture Notes in Comput. Sci., pages 355–370. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

Desharnais:2003:ALM

[DGJP03] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Approximating labelled Markov processes. Information and Computation, 184(1):160– 200, 2003. CODEN INFCEC. ISSN 0890-5401 (print), 1090-2651 (electronic).

Desharnais:2004:MLM

[DGJP04] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Metrics for labelled Markov processes. Theoretical Computer Science, 318(3):323–354, 2004. CODEN TCSDIQ. ISSN 0304-3975 (print), 1879-2294 (electronic).

Desharnais:2010:WBS

[DGJP10] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Weak bisimulation is sound and complete for pCTL. Information and Computation, 208 (2):203–219, 2010. CODEN IN-FCEC. ISSN 0890-5401 (print), 1090-2651 (electronic).

Desharnais:2000:ALM

[DJGP00] Josée Desharnais, Radha Jagadeesan, Vineet Gupta, and Prakash Panangaden. Approximating labeled Markov processes. In 15th Annual IEEE Symposium on Logic in Computer Science (Santa Barbara, CA, 2000), pages 95–106. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000.

Gupta:1997:RTA

[GHJ97] V. Gupta, T. A. Henzinger, and R. Jagadeesan. Robust timed automata. Lecture Notes in Computer Science, 1201:331–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godefroid:2001:ABM

[GHJ01] Patrice Godefroid, Michael Huth, and Radha Jagadeesan. Abstractionbased model checking using modal transition systems. In CONCUR 2001–concurrency theory (Aalborg), volume 2154 of Lecture Notes in Comput. Sci., pages 426–440. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Godefroid:2002:AAU

[GJ02] Patrice Godefroid and Radha Jagadeesan. Automatic abstraction using generalized model checking. In Computer aided verification, volume 2404 of Lecture Notes in Comput. Sci., pages 137–150. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

Gupta:2006:ARR

[GJP06] Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Approximate reasoning for real-time probabilistic processes. Log. Methods Comput. Sci., 2(1):1:4, 23, 2006. ISSN 1860-5974.

Gupta:1996:HCH

[GJS96a] V. Gupta, R. Jagadeesan, and V. Saraswat. Hybrid cc, hybrid automata, and program verification. Lecture Notes in Computer Science, 1066:52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:MCC

[GJS96b] Vineet Gupta, Radha Jagadeesan, and Vijay A. Saraswat. Models for concurrent constraint programming. In Montanari and Sassone [MS96], pages 66–83. ISBN 3-540-61604-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1996.

Gupta:1996:TCC

[GJS96c] Vineet Gupta, Radha Jagadeesan, and Vijay A. Saraswat. Truly concurrent constraint programming. In Montanari and Sassone [MS96], pages 373–388. ISBN 3-540-61604-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1996.

Gupta:1997:PCC

[GJS97] Vineet Gupta, Radha Jagadeesan, and Vijay Saraswat. Probabilistic concurrent constraint programming. Lecture Notes in Computer Science, 1243:243–257, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1998:CCC

[GJS98] V. Gupta, R. Jagadeesan, and V. A.

Saraswat. Computing with continuous change. Science of Computer Programming, 30(1–2):3–49, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

Gupta:2002:TCC

[GJS02] V. Gupta, R. Jagadeesan, and V. A. Saraswat. Truly concurrent constraint programming. Theoretical Computer Science, 278(1-2): 223–255, 2002. CODEN TCSDIQ. ISSN 0304-3975 (print), 1879-2294 (electronic). Mathematical foundations of programming semantics (Boulder, CO, 1996).

Gupta:1995:PHC

[GJSB95] V. Gupta, R. Jagadeesan, V. Saraswat, and D. G. Bobrow. Programming in hybrid constraint languages. Lecture Notes in Computer Science, 999: 226–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huth:2004:DER

[HJS04] Michael R. A. Huth, Radha Jagadeesan, and David A. Schmidt. A domain equation for refinement of partial systems. Math. Structures Comput. Sci., 14(4):469–505, 2004. ISSN 0960-1295.

Jagadeesan:1989:DLP

[Jag89] R. Jagadeesan. L-domains and lossless powerdomains. In Main [Mai90], pages 364–372. CODEN LNCSD9. ISBN 0-387-52790-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M38 1990. Proceedings of the Fifth International Conference on the Mathematical Foundations of Programming Semantics.

Jagadeesan:1990:DLP

[Jag90] Radhakrishnan Jagadeesan. Ldomains and lossless powerdomains. In Mathematical foundations of programming semantics (New Orleans, LA, 1989), volume 442 of Lecture Notes in Comput. Sci., pages 364– 372. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

Jagadeesan:1995:CTC

[JJ95] L. Jagadeesan and R. Jagadeesan. Causality and true concurrency: A dataflow analysis of the pi-calculus. In Alagar and Nivat [AN95], pages 277–?? CODEN LNCSD9. ISBN 3-540-60043-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .I5713 1995.

Jagadeesan:2006:PRB

[JJPR06] Radha Jagadeesan, Alan Jeffrey, Corin Pitcher, and James Riely. λ-RBAC: programming with rolebased access control. In Automata, languages and programming. Part II, volume 4052 of Lecture Notes in Comput. Sci., pages 456–467. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Jagadeesan:2008:RPR

[JJPR08] Radha Jagadeesan, Alan Jeffrey, Corin Pitcher, and James Riely. λ-RBAC: programming with rolebased access control [reprint of MR2375571]. Log. Methods Comput. Sci., 4(1):1:2, 1–24, 2008. ISSN 1860-5974. Special issue: Conference “International Colloquium on Automata, Languages and Programming 2006”.

Jagadeesan:2006:TPP

[JJR06] Radha Jagadeesan, Alan Jeffrey, and James Riely. Typed parametric polymorphism for aspects. Science of Computer Programming, 63(3): 267–296, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jagadeesan:2005:TCS

[JNS05] Radha Jagadeesan, Gopalan Nadathur, and Vijay Saraswat. Testing concurrent systems: an interpretation of intuitionistic logic. In FSTTCS 2005: Foundations of software technology and theoretical computer science, volume 3821 of Lecture Notes in Comput. Sci., pages 517–528. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005.

Jagadeesan:1990:DTM

[JP90] Radha Jagadeesan and Prakash Panangaden. A domain-theoretic model for a higher-order process calculus. In Paterson [Pat90], pages 181–194. CODEN LNCSD9. ISBN 3-540-52826-1 (Berlin), 0-387-52826-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.443. Preprint available as Cornell TR 89-1058.

Jagadeesan:1992:ASH

[JP92] R. Jagadeesan and K. Pingali. Abstract semantics for a higher-order functional language with logic variables. In ACM [ACM92], pages 355– 366. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. ACM order number 54990.

Jagadeesan:1991:FAS

[JPP91] R. Jagadeesan, P. Panangaden, and K. Pingali. A fully-abstract semantics for a first order functional language with logic variables. ACM Transactions on Programming Languages and Systems, 13(4), 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). Preliminary version appeared in the Proceedings of the 4th IEEE Symposium on Logic in Computer Science, June 1989.

Main:1990:MFP

[Mai90] M. Main, editor. Mathematical foundations of programming semantics: 5th international conference, Tulane University, New Orleans, Louisiana, USA, March 29–April 1, 1989: proceedings, volume 442 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNCSD9. ISBN 0-387-52790-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .M38 1990. Proceedings of the Fifth International Conference on the Mathematical Foundations of Programming Semantics.

Montanari:1996:CCT

[MS96] U. (Ugo) Montanari and Vladimiro Sassone, editors. CONCUR ’96: concurrency theory: 7th International Conference, Pisa, Italy, August 26–29, 1996: proceedings, volume 1119 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61604-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I53 1996.

Paterson:1990:ALP

[Pat90] Michael S. Paterson, editor. Automata, languages, and programming: 17th international colloquium, Warwick University, England, July 16–20, 1990: proceedings, volume 443 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNCSD9. ISBN 3-540-52826-1 (Berlin), 0-387-52826-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.443. ICALP (International Colloquium on Automata, Languages, and Programming) was sponsored by the European Association for Theoretical Computer Science.

Saraswat:2005:CCP

[SJ05] Vijay Saraswat and Radha Jagadeesan. Concurrent clustered programming (extended abstract). In CONCUR 2005—concurrency theory, volume 3653 of Lecture Notes in Comput. Sci., pages 353–367. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005.

Saraswat:1994:FTC

[SJG94a] Vijay Saraswat, Radha Jagadeesan, and Vineet Gupta. Foundations of timed concurrent constraint programming. In Proceedings, Ninth Annual IEEE Symposium on Logic in Computer Science, Paris, France, 4–7 July, 1994, pages 71– 80. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Saraswat:1994:PTC

[SJG94b] Vijay Saraswat, Radha Jagadeesan, and Vinheet Gupta. Programming in timed concurrent constraint languages. In B. Mayoh, E. Tyugu, and J. Penjaam, editors, Constraint Programming: Proceedings 1993 NATO ASI Parnu, Estonia, NATO Advanced Science Institute Series, pages 361–410. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57859-5.

Saraswat:1996:TDC

[SJG96] Vijay Saraswat, Radha Jagadeesan, and Vineet Gupta. Timed default concurrent constraint programming. Journal of Symbolic Computation, 22(5–6):475–520, 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Preliminary version appeared in the Proceedings of the 22nd Annual ACM SIGPLAN-SIGACT Symposium on REFERENCES 10

the Principles of Programming Languages, January 1995.